# EARN Fairness: Explaining, Asking, Reviewing, and Negotiating Artificial Intelligence Fairness Metrics Among Stakeholders

### v2

@@ 1. Introduction
However, there are three primary issues when applying fairness metrics to assess AI fairness. First, choosing the right metrics becomes problematic due to the lack of clear, unified guidelines for metric selection, especially in the absence of a one-size-fits-all metric and the insufficiency of using a single metric (Mehrabi et al., 2021; Madaio et al., 2022; Castelnovo et al., 2022; Haider et al., 2024). Second, AI fairness metric selection, often led by AI experts lacking stakeholder input, risks overly depending on these experts’ perspectives and disregards fairness as a societal value with diverse interpretations among stakeholders (Ashktorab et al., 2023; Verma and Rubin, 2018; Mehrabi et al., 2021; Hort et al., 2024). This could lead to AI models deemed technically ”fair” by AI experts but not accepted in practice because they fail to reflect stakeholders’ views on fairness (Whittaker et al., 2018; Lee et al., 2020). Previous works (Wong, 2020; Cheng et al., 2021; Zhang et al., 2020; Srivastava et al., 2019; Saxena et al., 2019) have begun to identify stakeholders’ metric preferences, for example, Cheng et al. (Cheng et al., 2021) developed an interface with an interview protocol to elicit each stakeholder’s fairness notions in child maltreatment risk management. However, the range of solicited metrics remains limited, and more work is needed to make fairness metrics more understandable to stakeholders without AI expertise, ensuring fairness is accessible to a broader audience. Last, even when personal fairness perspectives are collected from stakeholders, previous work has shown that different stakeholders often have different fairness goals (Madaio et al., 2022; Cheng et al., 2021; Nakao et al., 2023; Lee et al., 2017; Haider et al., 2024). Although there is a growing recognition of the importance of negotiating views on fairness and establishing consensus (Engelbutzeder et al., 2023; Lee et al., 2019a, b), there are knowledge gaps about when stakeholders’ fairness metric preferences collide, whether consensus can be reached on metric selection, what kind of consensus is reached, and how it is reached. Given these issues, it becomes crucial to not only capture the diverse personal preferences of stakeholders but also investigate fair approaches for negotiating differences and reaching a consensus between multiple stakeholders, ensuring AI fairness aligns with societal values.
@@ 2.1. AI Fairness Metrics
  • Demographic Parity (Kusner et al., 2017) (also known as Statistical Parity (Dwork et al., 2012)): Both protected and unprotected group members have the same probability of being classified in the positive category.
@@ 2.2. Soliciting Stakeholders’ Fairness Metric Preferences
AI fairness experts need to select from numerous metrics in a specific application context, yet there is no agreement on the most appropriate metrics in specific contexts (Hort et al., 2024; Dablain et al., 2022). This has motivated numerous studies to solicit stakeholders’ personal metric preferences (Wong, 2020; Cheng et al., 2021; Saxena et al., 2019; Haider et al., 2024; Ilvento, 2020)
@@
Once personal preferences for measuring fairness have been elicited, there is still a need to manage the differences in personal preferences. For instance, in loan scenarios, data scientists and domain experts have different perspectives on fairness (Nakao et al., 2023), with data scientists preferring group fairness and domain experts focusing mainly on individual fairness. Even within the same stakeholder types, people hold different metric preferences, for example, in child maltreatment risk management scenarios, some prefer equalized odds, while some prefer statistical parity (Cheng et al., 2021). Hence, balancing diverse fairness needs and seeking consensus among stakeholders is crucial for implementing fair AI successfully. Current research primarily focuses on seeking consensus on AI model building or predictive outcomes through participatory design (Zhang et al., 2023; Lee et al., 2019b) or co-design (Stumpf et al., 2021; Nakao et al., 2022) for AI fairness democratization. For example, in the context of goods division, Lee et al. (Lee et al., 2019a) reached distribution outcome consensus by allowing individual adjustments first, then group discussions for collective decision-making on the allocations. Furthermore, Lee et al. (Lee et al., 2019b) introduced the ”WeBuildAI” Framework to achieve a consensus-based algorithm driven by voting for fair food donation. While recent research advocates and explores negotiating different fairness notions among individuals, it has not translated into practical, quantifiable fairness metrics (Engelbutzeder et al., 2023). To the best of our knowledge, stakeholders negotiating the consensus on fairness metric selection has not been explicitly explored.
@@ 3.2.1. Design Goals
Our target users are individuals with no background in AI. To help users comprehend and compare different metrics, we draw inspiration from prior research (Nakao et al., 2023; Cheng et al., 2021; Stumpf et al., 2021), and propose four design goals: a.) Contextual Understanding: help users develop a contextual understanding of the model by analyzing AI model predictions case-by-case, b.) Model Transparency: explaining basic AI concepts, offering model explanations and model performance information to enhance users’ comprehension on AI model, c.) Diverse Explanation and Comparison: elucidating metrics in diverse styles and plain language, indicating when a metric shows the model to be unfair, presenting fairness metric results at the case level and providing intuitive comparisons across metrics, and d.) Fairness Adjustability: allowing users to question biases in original data and AI model predictions, and dynamically tracking their impact on model fairness. Moreover, our UI design and metric explanation methods are fundamentally rooted in the underlying mathematical logic of the metrics, allowing them to be adaptive and applicable across different domains.
@@ 3.2.2. Fairness Metrics Covered in FEE
Our framework also allows fairness threshold settings. Group-level metric results range from 0% to 100%, indicating increasing disparities among (sub)groups. Therefore, a threshold of 0% means absolute fairness and higher values signify higher unfairness tolerance. In contrast, individual fairness metrics range from 0% to 100%, indicating the degree of fairness in treating individuals. Thus, a threshold of 100% means absolute fairness and lower values signify less unfairness tolerance.
@@
Figure 3. This is another example of instance-level explanation views for group fairness metrics related to the protected feature of ”Age”. After users set the condition to ”Job” and click ”CHECK INDIVIDUAL INSTANCES” for Conditional Statistical Parity (in Figure 2 C1), they will see the corresponding instance-level explanation. Users can switch between conditions using radio buttons, such as ”Credit History”.
@@ 3.2.3. UI components
Data Exploration: This component allows users to explore datasets and AI model predictions (Figure 1 A). Each row represents a test instance, with horizontal scrolling to view all features. The last three columns are fixed, showing ground-truth labels (”Rated Credit” by experts), predictions (”Predicted Credit” by the AI model), and model confidence (’Probability’) (Nakao et al., 2023). Model confidence is displayed using a color-coded bar chart, and hovering reveals probability values for various predictions. Users can sort and filter instances by feature values to identify potential biases against certain groups or individuals, such as filtering instances based on gender and prediction results to observe how the model performs for different genders.
Model Exploration: This component lowers the barrier to understanding AI through visual explanations of model parameters and performance (Figure 1 B). In the ”Concept Explanation” view, users can check the machine learning workflow from training to prediction and click for concept explanations like protected features and model accuracy. In the ”Model Performance & Explanation” view, users see visual explanations of model performance: pie charts depicting overall model accuracy and class-specific accuracy, and bar charts showing distributions of positive and negative predictions for different features’ values (or bins for continuous features). By clicking ”Feature Weights”, users can inspect feature weights, identifying the importance of each feature and its positive (green bar) or negative (red bar) impact on prediction outcomes. This enables users to grasp which features are pivotal in predictions and assess if these features might lead to unfair outcomes.
@@ 3.2.4. Comparison with previous fairness UIs
Leveraging Nakao et al.’s work (Nakao et al., 2023) which provided user requirements of fairness investigation tools in a loan scenario  [^1] We extracted AI fairness exploration requirements from loan officers’ perspective, as we believe this group more closely aligns with our target audience, who are stakeholders without AI knowledge, Table 1 summarizes a comparison between our UI, FEE, and two other tools, FET (Cheng et al., 2021) and FairHIL (Nakao et al., 2023). Our tool, FEE, differs by supporting custom threshold inputs and dynamically explaining when metrics indicate fairness or unfairness. It also enhances the transparency of AI fairness by providing detailed model parameters and instance-level fairness metric explanations.
@@ 4.1. Data and Model
We applied the German Credit Dataset (Hofmann, 1994) due to its popularity in AI fairness research and providing a familiar high-risk banking scenario. It comprises 1000 instances, each with 20 attributes (13 categorical and 7 numerical features) for classifying ”Good” or ”Bad” credit, with no missing data. Our study investigated fairness on three features: age and gender, recognized as protected characteristics under law regulations  [^2] Equality Act 2010. https://www.legislation.gov.uk/ukpga/2010/15/contents, and foreign worker status as a sensitive feature utilized by many fairness research (Kamiran and Calders, 2009; Friedler et al., 2019). They were processed as follows 1) Age: we converted age into a binary feature, with age¡25 as the protected group; 2) Gender: we extracted this from the ”personal status and gender” attribute, and selected ”female” as the protected group; 3) Foreign worker: we chose ”yes” as the protected (sensitive) group. We provided four non-protected features as possible legitimate features in a credit rating scenario: job, savings, employment, and credit history (Verma and Rubin, 2018) for measuring Conditional Statistical Parity.
@@ 4.2. Participants
We recruited 18 participants for the study through public and university social media channels, all of whom are decision subjects impacted by AI decisions and have no AI background. Participants were aged 18 to 47 (mean = 28), with diverse genders, professions, nationalities, and ethnicities (Appendix A.2). Each participant received a £20 Amazon Voucher as compensation. We obtained informed consent from them, and each participant was assigned a unique ID for anonymity throughout the study. This study was reviewed and approved by the University of Glasgow Ethics Committee.
@@ 4.3. Procedure
Finally, we followed the framework to conduct team sessions for each team to complete Review and Negotiate. Members were provided with hard copies outlining their task and engaged in free-form negotiation without interference from researchers. Guidance was provided that the negotiation results should be unanimously approved by all team members. During the negotiation process, members were free to use FEE. Researchers aimed to ensure a harmonious negotiation atmosphere and offered alternative collective decision-making methods when needed, such as the Borda voting method due to its simplicity and robustness (Lee et al., 2019b; Kahng et al., 2019).
@@ 4.4. Data Collection and Analysis
We collected participants’ top 3 metric ranking lists and fairness thresholds, along with audio recordings and the researcher’s observation notes. We used descriptive statistics to analyze the top 3 metric ranking lists, reporting the raw frequencies for each metric within the top 3 ranks, and to calculate the mean values and variances for fairness thresholds. Further, we transcribed approximately 20 hours of recordings from all 18 participants. We conducted a thematic analysis (Braun and Clarke, 2006) to generate insights into participants’ diverse preferences on fairness metrics and thresholds, the reasons behind them, and user feedback on FEE design. Our research team met weekly to iterate over the themes and coding, reviewing and validating them. After 6 rounds of iteration, we finalized the final codebook (Appendix A.3).
@@ 5.1.1. Preferences for Fairness Metrics
We also observed that there is a wide variability of thresholds in Table 2, driven by a range of reasons (Appendix A.3.1). No two participants set exactly the same thresholds for all three fairness categories. In group and subgroup fairness settings, only two participants chose a 0% threshold, while four opted for a 100% threshold in individual fairness. We can note that on average participants set thresholds relatively higher than expected by law and by most practices (Gupta et al., 2021), with a group fairness mean of 9.28% (SD=7.51%), a subgroup fairness of 13.00% (SD=9.96%), and an individual fairness of 92.44% (SD=8.53%).
Most participants were willing to leave some room for unfairness, as it was deemed impossible to achieve in practice. As P14 remarked, ”We don’t need to be too absolute, … Experts or AI are not 100% accurate, that absolute perfection in fairness and accuracy is unattainable”. However, a small number of participants pursued perfect fairness, reflecting a zero-tolerance attitude towards bias. Another interesting observation is that participants tended to set higher thresholds for subgroup fairness, as more features were being considered compared to group fairness, thereby allowing for more leeway. As P10 stated, ”Since you’re taking more factors into account when assessing fairness, like more individual factors, you can maybe increase this a bit more, like […] more leeway”.
@@ 5.2. How do stakeholders negotiate to resolve personal differences in metric preferences? (RQ2)
In Team 3, reasons for and against fairness metrics were also discussed but they used a debate style, encouraging the presentation of arguments for and against certain fairness metrics. In the early negotiation stages, some polarization emerged similar to Team 1, but was successfully overcome through argumentation. For example, P10, P13, and P14 supported Counterfactual Fairness, while P9 and P18 favored Conditional Statistical Parity. In this team, we observed that certain team members dominated the discussions, leading to significant variations in the levels of participation among the group members. P10 presented arguments for Counterfactual Fairness while P18 supported Conditional Statistical Parity, leading two sides. The discussion primarily revolved around P18’s focus on group-level legitimate factors for fairness versus P10’s approach of directly excluding the impact of protected attributes for AI decision-making fairness, consulting FEE as needed to support their arguments. P10 proposed fairness as follows: ”It’s like oh no matter the value of age, gender, foreign worker is changed or not, AI prediction should not be changed […] You eliminate the bias first and then you think about employment [and other individual qualifications]”. After this perspective was accepted, Team 3, different from Team 1 and Team 2, advocated for using Counterfactual Fairness for individual fairness first, and then using Conditional Statistical Parity with a focus on subgroup fairness. Additionally, there was unanimous agreement on the importance of including at least age and gender to secure subgroup fairness in the Conditional Statistical Parity application.
Main finding 3: Teams used different strategies to come to a consensus. Voting was frequently employed but there were differences in how to decide the outcomes of voting. Reasons for or against metrics were discussed in all teams but this could also lead to entrenchments of opinions if not carefully handled. All teams chose to overcome conflict by combining metrics in some way, ending up with compromise solutions. The consensus showed that all teams considered both subgroup fairness and individual fairness, but with different priorities.

### v3

@@ Abstract.
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 2article: CSCW010publicationmonth: 4doi: 10.1145/3710908ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
However, there are three primary issues when applying fairness metrics to assess AI fairness. First, choosing the right metrics becomes problematic due to the lack of clear, unified guidelines for metric selection, especially in the absence of a ”one-size-fits-all” metric and the insufficiency of using a single metric (Mehrabi et al., 2021; Madaio et al., 2022; Castelnovo et al., 2022; Haider et al., 2024). Second, AI fairness metric selection, often led by AI experts lacking stakeholder input, risks overly depending on these experts’ perspectives and disregards fairness as a societal value with diverse interpretations among stakeholders (Ashktorab et al., 2023; Verma and Rubin, 2018; Mehrabi et al., 2021; Hort et al., 2024). This could lead to AI models deemed technically ”fair” by AI experts but not accepted in practice because they fail to reflect stakeholders’ views on fairness (Whittaker et al., 2018; Lee et al., 2020). Previous works (Wong, 2020; Cheng et al., 2021; Zhang et al., 2020; Srivastava et al., 2019; Saxena et al., 2019) have begun to identify stakeholders’ metric preferences, for example, Cheng et al. (Cheng et al., 2021) developed an interface with an interview protocol to elicit each stakeholder’s fairness notions in child maltreatment risk management. However, the range of solicited metrics remains limited, and more work is needed to make fairness metrics more understandable to stakeholders without AI expertise, ensuring fairness is accessible to a broader audience. Last, even when personal fairness perspectives are collected from stakeholders, previous work has shown that different stakeholders often have different fairness goals (Madaio et al., 2022; Cheng et al., 2021; Nakao et al., 2023; Lee et al., 2017; Haider et al., 2024). Although there is a growing recognition of the importance of negotiating views on fairness and establishing consensus (Engelbutzeder et al., 2023; Lee et al., 2019a, b), there are knowledge gaps about when stakeholders’ fairness metric preferences collide, whether consensus can be reached on metric selection, what kind of consensus is reached, and how it is reached. Given these issues, it becomes crucial to not only capture the diverse personal preferences of stakeholders but also investigate fair approaches for negotiating differences and reaching a consensus between multiple stakeholders, ensuring AI fairness aligns with societal values.
@@ 2.1. AI Fairness Metrics
  • Demographic Parity (Kusner et al., 2017) (also known as Statistical Parity (Dwork et al., 2012)): Both protected and unprotected group members have the same probability of being predicted as a positive class.
@@ 2.2. Soliciting Stakeholders’ Fairness Metric Preferences
AI fairness experts need to select from numerous metrics in a specific application context, yet there is no agreement on the most appropriate metrics in specific contexts (Hort et al., 2024; Dablain et al., 2022). This has motivated numerous studies to solicit stakeholders’ personal metric preferences (Wong, 2020; Cheng et al., 2021; Saxena et al., 2019; Haider et al., 2024; Ilvento, 2020).
@@
Once personal preferences for measuring fairness have been elicited, there is still a need to manage the differences in personal preferences. For instance, in loan scenarios, data scientists and domain experts have different perspectives on fairness (Nakao et al., 2023), with data scientists preferring group fairness and domain experts focusing mainly on individual fairness. Even within the same stakeholder types, people hold different metric preferences, for example, in child maltreatment risk management scenarios, some prefer equalized odds, while some prefer statistical parity (Cheng et al., 2021). Hence, balancing diverse fairness needs and seeking consensus among stakeholders is crucial for implementing fair AI successfully. Current research primarily focuses on seeking consensus on AI model building or predictive outcomes through participatory design (Zhang et al., 2023; Lee et al., 2019b) or co-design (Stumpf et al., 2021; Nakao et al., 2022) for AI fairness democratization. For example, in the context of goods division, Lee et al. (Lee et al., 2019a) reached distribution outcome consensus by allowing individual adjustments first, then group discussions for collective decision-making on the allocations. Furthermore, Lee et al. (Lee et al., 2019b) introduced the ”WeBuildAI” Framework to achieve a consensus-based algorithm driven by voting for fair food donation. While recent research advocates and explores negotiating different fairness notions among individuals, it has not translated into practical, quantifiable fairness metrics (Engelbutzeder et al., 2023). To the best of our knowledge, the process of stakeholders negotiating the consensus on fairness metric selection has not been explicitly explored.
@@ 3.2.1. Design Goals
Our target users are individuals with no background in AI. To help users comprehend and compare different metrics, we draw inspiration from prior research (Nakao et al., 2023; Cheng et al., 2021; Stumpf et al., 2021), and propose four design goals: a.) Contextual Understanding: help users develop a contextual understanding of the model by analyzing AI model predictions case-by-case, b.) Model Transparency: explaining basic AI concepts, offering model explanations and model performance information to enhance users’ comprehension of the AI model, c.) Diverse Explanation and Comparison: elucidating metrics in diverse styles and plain language, indicating when a metric shows the model to be unfair, presenting fairness metric results at the case level and providing intuitive comparisons across metrics, and d.) Fairness Adjustability: allowing users to question biases in original data and AI model predictions, and dynamically tracking their impact on model fairness. Moreover, our UI design and metric explanation methods are fundamentally rooted in the underlying mathematical logic of the metrics, allowing them to be adaptive and applicable across different domains.
@@ 3.2.2. Fairness Metrics Covered in FEE
Our framework also allows fairness threshold settings. Group-level metric results range from 0% to 100%, indicating increasing disparities among (sub)groups. Therefore, a threshold of 0% means absolute fairness and higher values signify higher unfairness tolerance. In contrast, individual fairness metric results range from 0% to 100%, indicating the degree of fairness in treating individuals. Thus, a threshold of 100% means absolute fairness and lower values signify higher unfairness tolerance.
@@
Figure 3. This is another example of instance-level explanation views for group fairness metrics related to the protected feature of ”Gender”. After users set the condition to ”Job” and click ”CHECK INDIVIDUAL INSTANCES” for Conditional Statistical Parity (in Figure 2 C1), they will see the corresponding instance-level explanation. Users can switch between conditions using radio buttons, such as ”Credit History”.
@@ 3.2.3. UI Components
Data Exploration: This component allows users to explore datasets and AI model predictions (Figure 1 A). Each row represents a test instance, with horizontal scrolling to view all features. The last three columns are fixed, showing ground-truth labels (”Rated Credit” by experts), predictions (”Predicted Credit” by the AI model), and model confidence (”Probability”) (Nakao et al., 2023). Model confidence is displayed using a color-coded bar chart, and hovering reveals probability values for various predictions. Users can sort and filter instances by feature values to identify potential biases against certain groups or individuals, such as filtering instances based on gender and prediction results to observe how the model performs for different genders.
Model Exploration: This component lowers the barrier to understanding AI through visual explanations of model parameters and performance (Figure 1 B). In the ”Concept Explanation” view, users can check the AI model’s workflow for prediction and click for concept explanations like protected features and model accuracy. In the ”Model Performance & Explanation” view, users see visual explanations of model performance: pie charts depicting overall model accuracy and class-specific accuracy, and bar charts showing distributions of positive and negative predictions for different features’ values (or bins for continuous features). By clicking ”MODEL WEIGHT”, users can inspect feature weights, identifying the importance of each feature and its positive (green bar) or negative (red bar) impact on prediction outcomes. This enables users to grasp which features are pivotal in predictions and assess if these features might lead to unfair outcomes.
@@ 3.2.4. Comparison with Previous Fairness UIs
Leveraging Nakao et al.’s work (Nakao et al., 2023) which provided user requirements of fairness investigation tools in a loan scenario  [^1] We extracted AI fairness exploration requirements from loan officers’ perspective, as we believe this group more closely aligns with our target audience, who are stakeholders without AI knowledge., Table 1 summarizes a comparison between our UI, FEE, and two other tools, FET (Cheng et al., 2021) and FairHIL (Nakao et al., 2023). Our tool, FEE, differs by supporting custom threshold inputs and dynamically explaining when metrics indicate fairness or unfairness. It also enhances the transparency of AI fairness by providing detailed model parameters and instance-level fairness metric explanations.
@@ 4.1. Data and Model
We applied the German Credit Dataset (Hofmann, 1994) due to its popularity in AI fairness research and providing a familiar high-risk banking scenario. It comprises 1000 instances, each with 20 features (13 categorical and 7 numerical features) for classifying ”Good” or ”Bad” credit, with no missing data. Our study investigated fairness on three features: age and gender, recognized as protected characteristics under law regulations  [^2] Equality Act 2010: https://www.legislation.gov.uk/ukpga/2010/15/contents
Charter of Fundamental Rights of the European Union, 2012: https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX:12012P/TXT
, and foreign worker status as a sensitive feature utilized by many fairness research (Kamiran and Calders, 2009; Friedler et al., 2019). They were processed as follows 1) Age: we converted age into a binary feature, with age¡25 as the protected group; 2) Gender: we extracted this from the ”personal status and gender” feature, and selected ”female” as the protected group; 3) Foreign Worker: we chose ”yes” as the protected (sensitive) group. We provided four non-protected features as possible legitimate features in a credit rating scenario: job, savings, employment, and credit history (Verma and Rubin, 2018) for measuring Conditional Statistical Parity.
@@ 4.2. Participants
We recruited 18 participants for the study through public and university social media channels, all of whom are decision subjects impacted by AI decisions and have no AI background. Participants were aged 18 to 47 (mean = 28), with diverse genders, professions, nationalities, and ethnicities (Appendix A.2). Each participant received a £20 Amazon voucher as compensation. We obtained informed consent from them, and each participant was assigned a unique ID for anonymity throughout the study. This study was reviewed and approved by the University of Glasgow College of Science & Engineering Ethics Committee (Application Number: 300230024).
@@ 4.3. Procedure
Finally, we followed the framework to conduct team sessions for each team to complete Review and Negotiate. Members were provided with hard copies outlining their tasks and engaged in free-form negotiation without interference from researchers. Guidance was provided that the negotiation results should be unanimously approved by all team members. During the negotiation process, members were free to use FEE. Researchers aimed to ensure a harmonious negotiation atmosphere and made preparations to offer alternative collective decision-making methods when needed, such as the Borda voting method due to its simplicity and robustness (Lee et al., 2019b; Kahng et al., 2019).
@@ 4.4. Data Collection and Analysis
We collected participants’ top 3 metric ranking lists and fairness thresholds, along with audio recordings and the researcher’s observation notes. We used descriptive statistics to analyze the top 3 metric ranking lists, reporting the raw frequencies for each metric within the top 3 ranks, and to calculate the mean values and standard deviations (SD) for fairness thresholds. Further, we transcribed approximately 20 hours of recordings from all 18 participants. We conducted a thematic analysis (Braun and Clarke, 2006) to generate insights into participants’ diverse preferences on fairness metrics and thresholds, the reasons behind them, and user feedback on FEE design. Our research team met weekly to iterate over the themes and coding, reviewing and validating them. After 6 rounds of iteration, we finalized the final codebook (Appendix A.3).
@@ 5.1.1. Preferences for Fairness Metrics
We also observed that there is a wide variability of thresholds in Table 2, driven by a range of reasons (Appendix A.3.1). No two participants set exactly the same thresholds for all three fairness categories. In group and subgroup fairness settings, only two participants chose a 0% threshold, while four opted for a 100% threshold in individual fairness. We can note that on average participants set thresholds relatively stricter than expected by law and by most practices (Gupta et al., 2021), with a group fairness mean of 9.28% (SD=7.51%), a subgroup fairness of 13.00% (SD=9.96%), and an individual fairness of 92.44% (SD=8.53%).
Most participants were willing to leave some room for unfairness, as it was deemed impossible to achieve in practice. As P14 remarked, ”We don’t need to be too absolute, … Experts or AI are not 100% accurate, that absolute perfection in fairness and accuracy is unattainable”. However, a small number of participants pursued perfect fairness, reflecting a zero-tolerance attitude towards bias. Another interesting observation is that participants tended to set more lenient thresholds for subgroup fairness, as more features being considered than in group fairness, thereby allowing for more leeway. As P10 stated, ”Since you’re taking more factors into account when assessing fairness, like more individual factors, you can maybe increase this a bit more, like […] more leeway”.
@@ 5.2. How do stakeholders negotiate to resolve personal differences in metric preferences? (RQ2)
In Team 3, reasons for and against fairness metrics were also discussed but they used a debate style, encouraging the presentation of arguments for and against certain fairness metrics. In the early negotiation stages, some polarization emerged similar to Team 1, but was successfully overcome through argumentation. For example, P10, P13, and P14 supported Counterfactual Fairness, while P9 and P18 favored Conditional Statistical Parity. In this team, we observed that certain team members dominated the discussions, leading to significant variations in the levels of participation among the group members. P10 presented arguments for Counterfactual Fairness while P18 supported Conditional Statistical Parity, leading two sides. The discussion primarily revolved around P18’s focus on group-level legitimate features for fairness versus P10’s approach of directly excluding the impact of protected features for AI decision-making fairness, consulting FEE as needed to support their arguments. P10 proposed fairness as follows: ”It’s like oh no matter the value of age, gender, foreign worker is changed or not, AI prediction should not be changed […] You eliminate the bias first and then you think about employment [and other individual qualifications]”. After this perspective was accepted, Team 3, different from Team 1 and Team 2, advocated for using Counterfactual Fairness for individual fairness first, and then using Conditional Statistical Parity with a focus on subgroup fairness. Additionally, there was unanimous agreement on the importance of including at least age and gender to secure subgroup fairness in the Conditional Statistical Parity application.
Main finding 3: Teams used different strategies to come to metric consensus. Voting was frequently employed but there were differences in how to decide the outcomes of voting. Reasons for or against metrics were discussed in all teams but this could also lead to entrenchments of opinions if not carefully handled. All teams chose to overcome conflict by combining metrics in some way, ending up with compromise solutions. The consensus showed that all teams considered both subgroup fairness and individual fairness, but with different priorities.